OpenStudy (anonymous):

New Question! (3x^1/3)-1=5

OpenStudy (anonymous):

Are you looking to solve for X?

OpenStudy (anonymous):

if solving for... x=8

OpenStudy (anonymous):

\[{3}^{1/3}-1=5\]

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

3 x^(1/3)-1 = 5 Isolate terms with x to the left hand side. Add 1 to both sides: 3 x^(1/3) = 6 Divide both sides by a constant to simplify the equation. Divide both sides by 3: x^(1/3) = 2 Eliminate the cube root on the left hand side. Raise both sides to the power of three: then u ll get x = 8

OpenStudy (anonymous):

I see. I suppose I just couldn't figure out how to cube root on my TI-84... I'll have to look that up

OpenStudy (anonymous):

^ (3)

OpenStudy (anonymous):

You will see the " ^ " button on the middle right side above the division symbol.

OpenStudy (anonymous):

right, but that won't help me cube root, will it?

OpenStudy (anonymous):

Cube yes but the cube root is actually " ^(1/3) "

OpenStudy (anonymous):

\[\sqrt[3]{x} = x ^{1/3}\]

OpenStudy (anonymous):

\[\sqrt[3]{x} = x^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

So you can use that button to solve for the cube root or any other root as well.

OpenStudy (anonymous):

Nice @Johnbc, how did you get it to do the x^1/3 like that?

OpenStudy (anonymous):

so, i'm down to x^1/3= 2, and I eliminate the ^1/3 how on my calculator?

OpenStudy (anonymous):

Last bar on the right side and last image as well. @Cosmichaotic

OpenStudy (anonymous):

Oh duh. Thank you much @Johnbc

OpenStudy (anonymous):

Well you need to perform algebra first and take the cube of both sides which is " ^ (3) "

OpenStudy (anonymous):

My pleasure.

OpenStudy (anonymous):